OpenStudy (anonymous):

u = <√7i+√8j>, v = <√8i-√7j>

OpenStudy (anonymous):

question?

OpenStudy (anonymous):

Find u x v and the angle between u and v to the nearest degree:

OpenStudy (anonymous):

your attempt?

OpenStudy (anonymous):

none... because the formulas online doesnt fit with my given question...

OpenStudy (anonymous):

what is the formula ?

OpenStudy (anonymous):

you mean the formula give you u =< , > and this problem give you u =< +...> right?

OpenStudy (anonymous):

just pick i, k and + out, you have the form of the formula

OpenStudy (anonymous):

I mean u =<sqrt7, sqrt8> v= <sqrt8, -sqrt 7> problem is solvable

OpenStudy (anonymous):

You should do it by yourself. It's not hard. You have formula, just APPLY. Dat sit. I can give you the answer in 1minute. But I wish you can handle it by YOURSELF

OpenStudy (anonymous):

i cant... bcoz im absent in that day... :(

OpenStudy (anonymous):

and i dont even know the title of that lesson...

OpenStudy (anonymous):

That is an unacceptable reason. You were absent, try to get the knowledge back is your duty. OK, if you want to study, I can help. If you don't, It's perfectly OK to me.

OpenStudy (anonymous):

ok.... just tell me the formula and i will do the rest... and im absent bcoz of measles...

OpenStudy (anonymous):

That was the past. Now, you are dealing with the problem you don't know how to solve. Be strong cos (theta) =

OpenStudy (anonymous):

numerator: u dot v denominator: magnitude of u * magnitude of v I am sorry for inconvenience, the draw box doesn't work

OpenStudy (anonymous):

thanks for encouragement

OpenStudy (anonymous):

u dot v = < u1*v1 + u2 *v2>

OpenStudy (anonymous):

u =<u1, u2> v = <v1,v2> so, u dot v = ???

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

56i^2-56j^2 ?

OpenStudy (anonymous):

take i and j out, nothing to do with them here, you are correct , =??

OpenStudy (anonymous):

wait lets try this 1... this is easier u = < –2, 9 > and v = < 3, –3 >

OpenStudy (anonymous):

one more thing , it is sqrt (56) -sqrt(56) =?? not 56 -56

OpenStudy (anonymous):

no,no, just continue

OpenStudy (anonymous):

a -a =?

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

you got the answer!!! check the option u dot v = 0 (one of the option)

OpenStudy (anonymous):

yeah the 3rd one ^^ tnx :)

OpenStudy (anonymous):

You see!! You CAN do it.

OpenStudy (anonymous):

u = < –2, 9 > and v = < 3, –3 > lets try this one :)

OpenStudy (anonymous):

ok, as usual, you have 2 parts

OpenStudy (anonymous):

cos (theta)= numerator: u dot v denominator: magnitude of u * magnitude of v solve numerator first

OpenStudy (anonymous):

uhm... numerator is 33? am i right ?

OpenStudy (anonymous):

-33, not 33 be careful with the sign

OpenStudy (anonymous):

huh? how come its -? <-2*-3> right ?

OpenStudy (anonymous):

no,not yet, you still have denominator: magnitude of u * magnitude of v u =<-2,9> ||u|| = sqrt((-2)^2 + 9^2)=?

OpenStudy (anonymous):

oh, sorry, your question is about numerator

OpenStudy (anonymous):

wait... we're still on numerator... <-2*-3>+<9*3>

OpenStudy (anonymous):

ok, I answer it now u = <-2 , 9> v = < 3, -3> u dot v = < (-2) *3 + 9* (-3)> = -6 - 27 = -33

OpenStudy (anonymous):

u1v1 + u2v2

OpenStudy (anonymous):

yeahyeah yeah ^^ my mistake sorry ^^ got confused sorry ^^,

OpenStudy (anonymous):

no worry. next the denominator!! I gave you instruction above. DO IT

OpenStudy (anonymous):

||u||*||v|| ?

OpenStudy (anonymous):

wait... what is the formula for magnitude?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

u = 85 v = 0 ?

OpenStudy (anonymous):

read my formula above ||u|| = sqrt ((u1)^2 +(u2)^2)

OpenStudy (anonymous):

square root, not just that

OpenStudy (anonymous):

ooops

OpenStudy (anonymous):

square root of (85)

OpenStudy (anonymous):

v is not =0, redo

OpenStudy (anonymous):

√-2^2+9^2 = √77 ?

OpenStudy (anonymous):

\[||u||=\sqrt{(-2)^2 +9^2}=\sqrt{4+81}=\sqrt{85}\]

OpenStudy (anonymous):

√3^2+3^2 = 3√2

OpenStudy (anonymous):

sqrt85 = 9.21 ?

OpenStudy (anonymous):

nope \[||v||=\sqrt{3^2 + (-3)^2}=\sqrt{9+9}=\sqrt{18}\]

OpenStudy (anonymous):

now, put all of them into the formula of cos \[cos (\theta )= \frac{u\dot v}{||u||*||v||=\frac{-33}{\sqrt{85}*\sqrt{18}}\] use calculator to find out the answer

OpenStudy (anonymous):

-0.84?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

again, use calculator to find theta by hitting cos^- of (-0.84)

OpenStudy (anonymous):

shift+cos?

OpenStudy (anonymous):

147.14?

OpenStudy (anonymous):

2nd + cos

OpenStudy (anonymous):

cos^-1(-0.84 = 147.1401196

OpenStudy (anonymous):

oh yea, you are right, hehehe

OpenStudy (anonymous):

hehehehe :"> tnx ^^youre a great help ^^ :*

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

^^